OpenStudy (anonymous):

How do you find the Cartesian equation for r=3sin(theta)?

Directrix (directrix):

For the conversion, three relationships may be needed. sin(theta) = ?/? and cos(theta) = ?/? and r^2 = x^2 + y^2 What goes in the ? fields for sin(theta) and cos(theta)?

Directrix (directrix):

@sharone Take a guess. Work with me.

OpenStudy (anonymous):

sin(theta)=1 cos(theta)=1 ? I know this is more than likely in correct...

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

incorrect*

Directrix (directrix):

sin(theta) = y/r and cos(theta) = x/r and then this: r^2 = x^2 + y^2

Directrix (directrix):

r=3sin(theta) so r = 3* (y/r) or ( (3y / r) ) r/1 = ( (3y / r) ) That will cross-multiply to give what result?

OpenStudy (anonymous):

@Directrix y=r^2/3

Directrix (directrix):

That's the same as r^2 = 3y. Right?

Directrix (directrix):

Now, we have to get rid of the r.

OpenStudy (anonymous):

would you replace the r with r=x/cos(theta)

Directrix (directrix):

No because to get to Cartesian, we can have only variables of x and y in this case. Polar has only r and theta so we have to get rid of those.

OpenStudy (anonymous):

Oh wait.. Right, okay..

Directrix (directrix):

r^2 = x^2 + y^2 --> Let's use this. Along with this: r^2 = 3y

OpenStudy (anonymous):

x^2+y^2=3y

OpenStudy (anonymous):

That would be the answer, right?

Directrix (directrix):

r^2 = x^2 + y^2 In place of r^2 in r^2 = 3y, substitute x^2 + y^2 to get x^2 + y^2 = 3y. Yes.

OpenStudy (anonymous):

Okay. Thank you so so much. You explained so much better than my teacher..

Directrix (directrix):

@sharone We might need to leave the equation in y = form. I don't know what is required. @matricked Thanks for your patience. Come on in to the discussion with us now.

OpenStudy (anonymous):

I can do the rest. Solving for the y and such. But thank you so much for your help!

Directrix (directrix):

This graph is a circle.